OpenStudy (anonymous):

Solve. Y + (-6) = -9 A) -15 B) -3 C) 54 D) 3

OpenStudy (anonymous):

move the -6 to the RHS.

OpenStudy (anonymous):

y-6=-9 y=-9+6 y=-3 so B

OpenStudy (anonymous):

What happens to the sign of -6 when you move it to the other side of the equal sign?

OpenStudy (anonymous):

@JayDS Good job. giving answers away just to get yourself a medal instead of allowing others to understand. Well done.

OpenStudy (raxis):

Oy, Jay... Look, if you add a negative, it's the same as subtracting, so Y plus negative 6 is the same as Y minus 6. You need to isolate Y, and you do this by doing the opposite with the numbers with it. Since you're subtracting 6 from Y, you add 6 to both sides to cancel 6 out and isolate Y. -9 plus 6.

OpenStudy (anonymous):

I'm showing him how to do it because I find it easier that way. Then if he is confused he can ask me questions, especially with these simple questions it's better looking at the working.

OpenStudy (anonymous):

@raxis, that's what I did....

OpenStudy (raxis):

I explained why you do it.

OpenStudy (anonymous):

kk.

OpenStudy (anonymous):

You shouldn't be giving answers away especially with these easy questions. That's when most people just take the answer instead of looking at the working.

OpenStudy (anonymous):

ok, I will keep that in mind @Azteck. sorry if I have caused any misunderstanding.

OpenStudy (anonymous):

Nah, it's just a heads up @JayDS

OpenStudy (anonymous):

kk, good to know.